The sum of two numbers is 50. Two times the smaller number exceeds the larger number by 19 . Find the numbers.
Answer by stanbon(75887)

The sum of two numbers is 50. Two times the smaller number exceeds the larger number by 19 . Find the numbers.
-----------------
x + y = 50
2x = y + 19
------
Rearrange:
2x + 2y = 100
2x - y = 19
-----
Subtract and solve for "y"::
3y = 81
y = 27
---
x = 50-27 = 23
-------------
Cheers,
Stan H.
-------------